What is the measure of a median of a trapezoid if the base is 15 and 18?



1
Expert's answer
2022-03-28T16:43:02-0400

The measure of a median of a trapezoid is calculated as:

"\begin{aligned} m = &\frac{bigger \ base +smaller \ base}{2} \\= &\frac{18 +15}{2}\\=& \frac{33}{2}\\ m = & 16.5\end{aligned}"